OpenStudy (anonymous):

Really need help! Can someone help me simplify a radical? I have this question: Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. But everything I come up with turns out weird. Can someone help me come up with an equation? I can try to solve it, I just need help figuring out what I need to do.

OpenStudy (anonymous):

here is an idea out of this problem X^4+3x^3-5x^2-21x+22=0 one of the solutions is -3+sprt2i is this what you need?

OpenStudy (anonymous):

I suppose I have to start out with something like √98x^5 and then simplify from there. Is that a simplified version of something?

OpenStudy (mertsj):

Let's just follow the directions. "The radicand consists of a coefficient and a variable. That could be something simple like 5x. But wait...it says the coefficient should be simplified using the greatest perfect square method. How about 20? So let's just use 20x for our radicand: \[\sqrt{20x}\]

OpenStudy (mertsj):

Now it says to simplify it completely: \[\sqrt{20x}=\sqrt{4(5)x}=2\sqrt{5x}\]

OpenStudy (mertsj):

And we're done

OpenStudy (anonymous):

Oh wow, thank you so much!

OpenStudy (mertsj):

yw